OpenStudy (goformit100):

The ages of Hari and Harry are in the ratio 5:7. Four years from now the ratio of their ages will be 3:4. Find their present ages.

OpenStudy (goformit100):

@Justin9529

OpenStudy (justin9529):

Hari = 5x years old Harry = 7x years old (5x +4)/(7x +4) = 3/4 [cross multiply] 4(5x +4) = 3(7x +4) 20x +16 = 21x +12 16 -12 = 21x - 20x 4 = x Hari = 5x years old = 5(4) = 20 years old Harry = 7x years old = 7(4) = 28 years old Check: (20+4)/(28+4) = 24/32 = 3/4 or 3:4

OpenStudy (goformit100):

THANK you Sir